# Poisson approximation of prime divisors of shifted primes

### v1

@@ 1. Introduction
The prime factorization of a random integer $n\leqslant x$ is well-approximated, in a probabilistic sense, by a sequence of independent random variables. This approximation is known as the Kubilius model after the pioneering work of Kubilius [30]; see also [7] for a development of the model. In this note we develop an analogous model for the prime factorization of a shifted prime $p+a$ with $a$ fixed and $p$ a prime drawn at random from $(|a|,x]$. We prove a total variation distance estimate for the difference between the model and actual prime factors of shifted primes, and apply it to show that the prime factors of shifted primes in disjoint sets behave like independent Poisson variables. This in turn is used to establish a general transference principle, which states roughly that any property of the prime factors of a random integer $n\leqslant x$, which is weakly dependent on the smallest and largest prime factors of $n$, also holds for the prime factors of a random shifted prime $p+a$ with $p\leqslant x$. This excludes properties such as bounding the size of the largest prime factor of $n$ (a notoriously difficult problem) and the property “$2|n$”, which must be excluded since it occurs for half of integers $n\leqslant x$, but for $\pi(x)-1$ primes $p\leqslant x$ with $n=p+a$ and $a$ is odd, and for only one prime $p$ when $a$ is even,
When $a\in\{-1,1\}$, the distribution of the prime factors of numbers $p+a$, $p$ being prime, plays a central role in investigations of Euler’s totient function and the sum of divisors function (see, for example, [9, 11, 10, 13, 14, 17, 20, 22, 35]), the Carmichael $\lambda$-function [12, 21], orders and primitive roots modulo primes [18, 23, 28, 29], Carmichael numbers [2, 25], and in computational problems (e.g. primality testing, factorization, pseudo-random number generators), see e.g., [2], [32] and Sections 3.3, 3.5, 4.1, 4.5 and 5.4 of [5].
@@ 1.1. Kubilius’ model of prime factors of integers
(1.1) $d_{TV}(X,Y):=\sup_{A\subset\Omega}\big{|}\mathbb{P}(X\in A)-\mathbb{P}(Y\in A)\big{|}.$
@@ 1.2. A Kubilius type model for shifted primes
Fix a nonzero integer $a$. It is expected that the distribution of the prime factors of a random shifted prime $p+a$ with $p\leqslant x$ behaves very much like the distribution of the prime factors of a random integer in $[1,x]$, provided that the statistic under consideration depends weakly on the smallest prime factors of $p+a$; this restriction is needed since if $q$ is prime and $q\nmid a$, then $q|(p+a)$ for a proportion $\frac{1}{q-1}$ of the primes by the prime number theorem for arithmetic progressions [6, §20]. This principle has been established for many statistics which are weakly dependent on both the very smallest prime factors and the very largest prime factors. The first such result is the seminal 1935 paper [9] of Erdős, who showed that the number of prime factors of $p+a$ is usually very close to $\log\log x$. This established an analog for shitfed primes of the famous theorem of Hardy and Ramanujan, who showed in 1917 that most integers $n\leqslant x$ has about $\log\log x$ prime divisors. To date the strongest results about the number of prime factors of $p+a$ are due to Timofeev [41] and the author [16]. The theory of general additive functions on shifted primes has been much studied, e.g. Elliott [8]. However, to the author’s knowledge nobody has written down an explicit comparison analogous to Theorem A.
@@
Also, the variables $W_{q}$ are mutually independent. Thus, by the prime number theorem for arithmetic progressions, $\mathbb{P}(W_{q}=v)$ is the density of primes $p$ for which $q^{v}\|(p+a)$. Let $\mathbf{W}_{y}$ be the vector of $W_{q}$ for $q\leqslant y$. For a randomly drawn prime $p\in(|a|,x]$ write
@@
and let $\mathbf{U}_{x,y}$ denote the vector $(u_{q}:q\leqslant y)$.
Our results depend on the “level of distribution” of primes in progressions with smooth moduli; Denote by $P^{+}(n)$ the largest prime factor of $n$, with the convention that $P^{+}(1)=0$. Let $P^{-}(n)$ be the smallest prime factor of $n$, with the convention that $P^{-}(1)=\infty$. As is usual, denote by $\pi(x)$ the number of primes $p\leqslant x$ and by $\pi(x;m,b)$ the number of primes $p\leqslant x$ in the progression $b\mod m$. We need that $\pi(x;m,b)$ is approximately $\frac{\pi(x)}{\phi(m)}$ uniformly on average over smooth moduli $m$ up to some power of $x$. Consider the following hypothesis.
@@
Hypothesis $Z(\frac{1}{2})$ (where we may take $\delta=1$ for any $\varepsilon>0$) is a consequence of the Bombieri-Vinogradov Theorem [6, Chapter 28], improving upon an earlier result of Barban [4], who showed Hypothesis $Z(\frac{3}{8})$, again with $\delta=1$ for any $\varepsilon>0$. That Hypothesis $Z(\gamma)$ holds for some $\gamma>1/2$ is due to Yitang Zhang [42] and was a crucial ingredient in his celebrated proof that the gaps between primes are infinitely often bounded by $7\cdot 10^{7}$. The subsequent PolyMath8a project [36] established Hypothesis $Z(\frac{1}{2}+\frac{7}{300})$, Maynard [34, Corollary 1.2] showed Hypothesis $Z(\frac{1}{2}+\frac{1}{42})$, and Stadlmann [38] proved Hypothesis $Z(\frac{1}{2}+\frac{1}{40})$. It is conjectured that $Z(1)$ holds; indeed, this is a simple consequence of the Elliott-Halberstam conjecture.
@@ 1.3. Poisson approximation of prime factors of shifted primes
The analogous statement for the distribution of prime factors of all integers $n\leqslant x$ has been given by the author [15]. Provided that the right side in the conclusion is sufficiently small, this reduces the problem of the distribution of $(f_{1},\ldots,f_{m})$ to the distribution of $(Z_{1},\ldots,Z_{m})$, the latter being much easier especially due to the independence of $Z_{1},\ldots,Z_{m}$.
@@
The first probability is over a random integer $n\leqslant x$ and the second is over a random prime $p$ satisfying $|a|<p\leqslant x$.
@@
Assume the hypotheses of Theorem 4, that $y\leqslant x^{o(1)}$ (as $x\to\infty$) and either (i) $\min_{j}H_{1}(T_{j})\to\infty$ as $x\to\infty$ or (ii) $\min_{j}H_{2}(T_{j})=o(1)$ as $x\to\infty$. Then, uniformly over all $\mathscr{R}\subset{\mathbb{N}}_{0}^{m}$ we have
$\mathbb{P}\big{(}(f_{1}(n),\ldots,f_{m}(n))\in\mathscr{R}\big{)}-\mathbb{P}\big{(}(f_{1}(p+a),\ldots,f_{m}(p+a))\in\mathscr{R}\big{)}\Big{|}=o(1)\qquad(x\to\infty).$
@@ 1.4. An application
$\sum_{j=2}^{k+1}\frac{H_{2}(T_{j})}{1+\sqrt{H_{1}(T_{j})}}+\mathrm{e}^{-\alpha u\log u}+\frac{1}{\log x}\ll\frac{1}{\log_{2}x}+\frac{1}{\sqrt{H_{1}(T_{1})}}+\sum_{q>\xi(x)}\frac{1}{q^{2}}\ll\frac{1}{\sqrt{\log_{2}\xi(x)}}.$
Therefore, (1.8) holds with $n=p+a$ for all but $o(x/\log x)$ primes $p\in(|a|,x]$, as $x\to\infty$. Recalling (1.7), this completes the proof. ∎
@@
Acknowledgements. The author thanks Steve Fan, Mikhail Gabdullin and Gérald Tenenbaum for helpful comments on an earlier version of the paper. The author was supported in part by National Science foundation grants DMS-1802139 and DMS-2301264.
@@ 2. Total variation distance inequalities
valid for all real numbers $a_{1},b_{1},\ldots,a_{m},b_{m}$. Denoting $\Omega$ the domain of $(X_{1},\ldots,X_{m})$, and writing $a_{i}=\mathbb{P}(X_{i}=\omega_{i})$, $b_{i}=\mathbb{P}(Y_{i}=\omega_{i})$, we then have
@@ 3. Sieving by small primes
$\sum_{m\in\mathcal{E}(x,y;\theta,A)}\frac{1}{\phi(m)}\ll_{A,\theta}\frac{1}{(\log x)^{A+1}}.$
@@
$\Phi_{m}(x,y)=\#\bigg{\{}|a|<p\leqslant x:p\equiv-a\allowbreak\mkern 7.0mu({\operator@font mod}\,\,m),P^{-}\left(\frac{p+a}{m}\right)>y\bigg{\}}.$
@@
[…]
Since #$\#\{b\in\mathcal{A}:q|b\}=\pi(x;qm,-a)$ for primes $q$, heuristically we expect $\Phi_{m}(x,y)\approx\pi(x)g_{y}(m)$.
We note that Mertens’ theorem implies that
@@
$\Phi_{m}(x,y)=\pi(x)g_{y}(m)\left(1+O_{\theta,\varepsilon}\left(e^{-(1-\varepsilon){u}_{m}^{\phantom{2}}\log({u}_{m}^{\phantom{2}}+1)}+\frac{1}{(\log x)^{A}}\right)\right),$
@@
(3.5) $\Phi_{m}(x,y)=\pi(x)g_{m}(y)\Big{(}1+O_{\varepsilon}(e^{-(1-\varepsilon)v\log v})\Big{)}+O\bigg{(}\sum_{\begin{subarray}{c}d\leqslant D\\ P^{+}(d)\leqslant y\\ (dm,a)=1\end{subarray}}\Big{|}\pi(x;dm,-a)-\frac{\pi(x)}{\phi(dm)}\Big{|}\bigg{)}$
@@
Taking $D=\max(y,x^{\theta}/m)$, so that $v=\max(1,{u}_{m}^{\phantom{2}})$. Thus, $v\log v={u}_{m}^{\phantom{2}}\log({u}_{m}^{\phantom{2}}+1)+O(1)$. Since $m\not\in\mathcal{E}(x,y;\theta,A+1)$, (3.4) implies that
$\sum_{\begin{subarray}{c}d\leqslant D\\ P^{+}(d)\leqslant y\\ (dm,a)=1\end{subarray}}\Big{|}\pi(x;dm,-a)-\frac{\pi(x)}{\phi(dm)}\Big{|}\leqslant\frac{\pi(x)}{\phi(m)(\log x)^{A+1}}\ll_{a}\frac{\pi(x)g_{y}(m)}{(\log x)^{A}}.\qed$
@@
If $y\geqslant 2$ and $t\geqslant 2$ then for any $\varepsilon>0$,
$\sum_{\begin{subarray}{c}m\geqslant t\\ P^{+}(m)\leqslant y\end{subarray}}\frac{1}{\phi(m)}\ll_{\varepsilon}(\log y)e^{-s\log(s+1)}+t^{-1+\varepsilon},\quad s=\frac{\log t}{\log y}.$
@@
Completely standard, thus we only sketch the proof. Assume $\varepsilon$ is sufficiently small. It suffices to prove the estimate in the case $(\log t)^{1+\varepsilon}\leqslant y\leqslant t^{\varepsilon}$, since the bound $O(\log y)$ is trivial (use it for $y>t^{\varepsilon}$) and for $y\leqslant(\log t)^{1+2\varepsilon}$, $u\log u\geqslant(1-2\varepsilon)\log t$ and thus the statement for $y<(\log t)^{1+2\varepsilon}$ follows from the bound for $y=(\log t)^{1+2\varepsilon}$. First write
@@
$\sum_{\begin{subarray}{c}m\geqslant t\\ P^{+}(m)\leqslant y\end{subarray}}\frac{1}{\phi(m)}=\sum_{P^{+}(d)\leqslant y}\frac{\mu^{2}(d)}{d\phi(d)}\sum_{\begin{subarray}{c}k\geqslant t/d\\ P^{+}(k)\leqslant y\end{subarray}}\frac{1}{k}.$
The inner sum is $O(\log y)$ and therefore the contribution from $d>t^{1-\varepsilon/2}$ is $O(t^{-1+\varepsilon})$. Using standard bounds on counts of numbers with only small prime factors, cf. [40, Ch. III.5], we readily obtain by partial summation
$\sum_{\begin{subarray}{c}k\geqslant t/d\\ P^{+}(k)\leqslant y\end{subarray}}\frac{1}{k}\ll_{\varepsilon}(\log y)e^{-v\log(v+1)}+\left(\frac{d}{t}\right)^{1-\varepsilon},$
where
$v=\frac{\log\frac{t}{d}}{\log y}=s-\frac{\log d}{\log y}.$
We note that $\frac{\log(s+1)+1}{\log y}\leqslant\frac{1}{1+\varepsilon}$ and thus by the mean value theorem,
$v\log(v+1)\geqslant s\log(s+1)-(\log(s+1)+1)\frac{\log d}{\log y}\geqslant s\log(s+1)-(1-\varepsilon)\log d.$
Summing on $d\leqslant t^{1-\varepsilon/2}$ yields the desired bound. ∎

### v3

@@ 1. Introduction
The prime factorization of a random integer $n\leqslant x$ is well-approximated, in a probabilistic sense, by a sequence of independent random variables. This approximation is known as the Kubilius model after the pioneering work of Kubilius [30]; see also [7] for a development of the model. In this note we develop an analogous model for the prime factorization of a shifted prime $p+a$ with $a$ fixed and $p$ a prime drawn at random from $(|a|,x]$. We prove a total variation distance estimate for the difference between the model and actual prime factors of shifted primes, and apply it to show that the prime factors of shifted primes in disjoint sets behave like independent Poisson variables. This in turn is used to establish a general transference principle, which states roughly that any property of the prime factors of a random integer $n\leqslant x$, which is weakly dependent on the smallest and largest prime factors of $n$, also holds for the prime factors of a random shifted prime $p+a$ with $p\leqslant x$. This excludes properties such as bounding the size of the largest prime factor of $n$ (a notoriously difficult problem) and the property “$2|n$”, which must be excluded since it occurs for half of integers $n\leqslant x$, but for $\pi(x)-1$ primes $p\leqslant x$ with $n=p+a$ when $a$ is odd, and for only one prime $p$ when $a$ is even.
When $a\in\{-1,1\}$, the distribution of the prime factors of numbers $p+a$, $p$ being prime, plays a central role in investigations of Euler’s totient function and the sum of divisors function (see, for example, [9, 10, 11, 13, 14, 17, 20, 22, 35]), the Carmichael $\lambda$-function [12, 21], orders and primitive roots modulo primes [18, 23, 28, 29], Carmichael numbers [2, 25], and in computational problems (e.g. primality testing, factorization, pseudo-random number generators), see e.g., [1], [2], [32] and Sections 3.3, 3.5, 4.1, 4.5 and 5.4 of [5].
@@ 1.1. Kubilius’ model of prime factors of integers
(1.1) $d_{TV}(X,Y):=\sup_{U\subset\Omega}\big{|}\mathbb{P}(X\in A)-\mathbb{P}(Y\in A)\big{|}.$
@@ 1.2. A Kubilius type model for shifted primes
Fix a nonzero integer $a$. It is expected that the distribution of the prime factors of a random shifted prime $p+a$ with $p\leqslant x$ behaves very much like the distribution of the prime factors of a random integer in $[1,x]$, provided that the statistic under consideration depends weakly on the smallest prime factors of $p+a$; this restriction is needed since if $q$ is prime and $q\nmid a$, then $q|(p+a)$ for a proportion $\frac{1}{q-1}$ of the primes by the prime number theorem for arithmetic progressions [6, §20]. This principle has been established for many statistics which are weakly dependent on both the very smallest prime factors and the very largest prime factors. The first such result is the seminal 1935 paper [9] of Erdős, who showed that the number of prime factors of $p+a$ is usually very close to $\log\log x$. This established an analog for shifted primes of the famous theorem of Hardy and Ramanujan, who showed in 1917 that most integers $n\leqslant x$ have about $\log\log x$ prime divisors. To date the strongest results about the number of prime factors of $p+a$ are due to Timofeev [41] and the author [16]. The theory of general additive functions on shifted primes has been much studied, e.g. Elliott [8]. However, to the author’s knowledge nobody has written down an explicit comparison analogous to Theorem A.
@@
Also, the variables $W_{q}$ are mutually independent. Thus, by the prime number theorem for arithmetic progressions, $\mathbb{P}(W_{q}=v)$ is the density of primes $p$ for which $q^{v}\|(p+a)$. Let $\mathbf{W}_{y}$ be the vector of $W_{q}$ for $q\leqslant y$. For a randomly drawn prime $p\in(|a|+1,x]$ write
@@
and let $\mathbf{U}_{x,y}$ denote the vector $(u_{q}:q\leqslant y)$. The lower bound on $p$ ensures that $p$ is odd and $p+a\geqslant 2$.
Our results depend on the “level of distribution” of primes in progressions with smooth moduli. Denote by $P^{+}(n)$ the largest prime factor of $n$, with the convention that $P^{+}(1)=0$. Let $P^{-}(n)$ be the smallest prime factor of $n$, with the convention that $P^{-}(1)=\infty$. As is usual, denote by $\pi(x)$ the number of primes $p\leqslant x$ and by $\pi(x;m,b)$ the number of primes $p\leqslant x$ in the progression $b\mod m$. We need that $\pi(x;m,b)$ is approximately $\frac{\pi(x)}{\phi(m)}$ uniformly on average over smooth moduli $m$ up to some power of $x$. Consider the following hypothesis.
@@
Hypothesis $Z(\frac{1}{2})$ (where we may take $\delta=1$ for any $\varepsilon>0$) is a consequence of the Bombieri-Vinogradov Theorem [6, Chapter 28], improving upon an earlier result of Barban [4], who showed Hypothesis $Z(\frac{3}{8})$, again with $\delta=1$ for any $\varepsilon>0$. The best known result is $Z(\frac{1}{2}+\frac{1}{42})$, due to Maynard [34, Corollary 1.2], which uses ideas originating from the celebrated work of Zhang [42] and further developed by the PolyMath8a project [36]. We note that the estimates of Zhang, PolyMath8b as well as new work of Stadlmann [38] establish versions of Hypothesis $Z(\gamma)$ with $\gamma>\frac{1}{2}$ but with the sum restricted to squarefree $m$. It is conjectured that $Z(1)$ holds; indeed, this is a simple consequence of the Elliott-Halberstam conjecture.
@@ 1.3. Poisson approximation of prime factors of shifted primes
The analogous statement for the distribution of prime factors of all integers $n\leqslant x$ has been given by the author [15]. Provided that the right side in the conclusion is sufficiently small, this reduces problems about the distribution of $(f_{1},\ldots,f_{m})$ to the analogous problems about the distribution of $(Z_{1},\ldots,Z_{m})$, the latter being much easier especially due to the independence of $Z_{1},\ldots,Z_{m}$.
@@
The first probability is over a random integer $n\leqslant x$ and the second is over a random prime $p$ satisfying $|a|+1<p\leqslant x$.
@@
Assume the hypotheses of Theorem 4, that $y\leqslant x^{o(1)}$ (as $x\to\infty$) and either (i) $\min_{j}H_{1}(T_{j})\to\infty$ as $x\to\infty$ or (ii) $\max_{j}H_{2}(T_{j})=o(1)$ as $x\to\infty$. Then, uniformly over all $\mathscr{R}\subset{\mathbb{N}}_{0}^{m}$ we have
$\Big{|}\mathbb{P}\big{(}(f_{1}(n),\ldots,f_{m}(n))\in\mathscr{R}\big{)}-\mathbb{P}\big{(}(f_{1}(p+a),\ldots,f_{m}(p+a))\in\mathscr{R}\big{)}\Big{|}=o(1)\qquad(x\to\infty).$
@@ 1.4. An application
$\sum_{j=1}^{k+1}\frac{H_{2}(T_{j})}{1+\sqrt{H_{1}(T_{j})}}+\mathrm{e}^{-\alpha u\log u}+\frac{1}{\log x}\ll\frac{1}{\log_{2}x}+\frac{1}{\sqrt{H_{1}(T_{1})}}+\sum_{q>\xi(x)}\frac{1}{q^{2}}\ll\frac{1}{\sqrt{\log_{2}\xi(x)}}.$
Therefore, (1.8) holds with $n=p+a$ for all but $o(x/\log x)$ primes $p\in(|a|+1,x]$, as $x\to\infty$. Recalling (1.7), this completes the proof. ∎
@@
Acknowledgements. The author thanks Steve Fan, Mikhail Gabdullin, Gérald Tenenbaum and the anonymous referees for helpful comments on earlier versions of the paper. The author was supported in part by National Science foundation grants DMS-1802139 and DMS-2301264.
@@ 2. Total variation distance inequalities
valid for all real numbers $a_{1},b_{1},\ldots,a_{m},b_{m}$. Denoting by $\Omega$ the domain of $(X_{1},\ldots,X_{m})$, and writing $a_{i}=\mathbb{P}(X_{i}=\omega_{i})$, $b_{i}=\mathbb{P}(Y_{i}=\omega_{i})$, we then have
@@ 3. Sieving by small primes
$\sum_{m\in\mathcal{E}(x,y;\theta,A)}\frac{1}{\phi(m)}\ll_{A,\theta,a}\,\frac{1}{(\log x)^{A+1}}.$
@@
$\Phi_{m}(x,y)=\#\bigg{\{}|a|+1<p\leqslant x:p\equiv-a\allowbreak\mkern 7.0mu({\operator@font mod}\,\,m),P^{-}\left(\frac{p+a}{m}\right)>y\bigg{\}}.$
@@
$\mathcal{A}=\Big{\{}\frac{p+a}{m}:|a|+1<p\leqslant x,p\equiv-a\allowbreak\mkern 7.0mu({\operator@font mod}\,\,m)\Big{\}}.$
Since #$\#\{b\in\mathcal{A}:q|b\}=\pi(x;qm,-a)+O_{a}(1)$ for primes $q$, heuristically we expect that $\Phi_{m}(x,y)$ is approximately $\pi(x)g_{y}(m)$.
We note that Mertens’ theorem and (3.3) imply that
@@
$\Phi_{m}(x,y)=\pi(x)g_{y}(m)\left(1+O_{a,\varepsilon}\left(e^{-(1-\varepsilon){u}_{m}^{\phantom{2}}\log({u}_{m}^{\phantom{2}}+1)}+\frac{1}{(\log x)^{A}}\right)\right),$
@@
(3.5) $\Phi_{m}(x,y)=\pi(x)g_{y}(m)\Big{(}1+O_{\varepsilon}(e^{-(1-\varepsilon)v\log v})\Big{)}+O_{a}\bigg{(}D+\sum_{\begin{subarray}{c}d\leqslant D\\ P^{+}(d)\leqslant y\\ (dm,a)=1\end{subarray}}\Big{|}\pi(x;dm,-a)-\frac{\pi(x)}{\phi(dm)}\Big{|}\bigg{)}$
@@
We take $D=\max(y,x^{\theta}/m)$, so that $v=\max(1,{u}_{m}^{\phantom{2}})$. Thus, $v\log v={u}_{m}^{\phantom{2}}\log({u}_{m}^{\phantom{2}}+1)+O(1)$. Since $m\not\in\mathcal{E}(x,y;\theta,A+1)$, (3.4) implies that
$\sum_{\begin{subarray}{c}d\leqslant D\\ P^{+}(d)\leqslant y\\ (dm,a)=1\end{subarray}}\Big{|}\pi(x;dm,-a)-\frac{\pi(x)}{\phi(dm)}\Big{|}\leqslant\frac{\pi(x)}{\phi(m)(\log x)^{A+1}}\ll_{a}\frac{\pi(x)g_{y}(m)}{(\log x)^{A}}.$
Finally, by (3.2), $D\ll\pi(x)g_{y}(m)(\log x)^{-A}$ as well. ∎
@@
If $y\geqslant 2$, $t\geqslant 2$ and $y\geqslant(\log t)^{2}$ then
$\sum_{\begin{subarray}{c}m\geqslant t\\ P^{+}(m)\leqslant y\end{subarray}}\frac{1}{\phi(m)}\ll(\log y)e^{-s\log(s+1)},\quad s=\frac{\log t}{\log y}.$
@@
Let $c>0$ be sufficiently large. The sum is trivially $O(\log y)$, and this suffices for $s\leqslant c$. Now assume $s>c$, so that $y<t^{1/c}$ and also $t\geqslant 2^{c}$; that is, we may assume that $t$ is sufficiently large. For $(\log t)^{2}\leqslant y\leqslant t^{1/c}$, write
@@
(3.6) $\sum_{\begin{subarray}{c}m\geqslant t\\ P^{+}(m)\leqslant y\end{subarray}}\frac{1}{\phi(m)}=\sum_{P^{+}(d)\leqslant y}\frac{\mu^{2}(d)}{d\phi(d)}\sum_{\begin{subarray}{c}k\geqslant t/d\\ P^{+}(k)\leqslant y\end{subarray}}\frac{1}{k}.$
The inner sum is $O(\log y)$ and therefore the contribution from $d>t^{2/3}$ is $O(t^{-2/3}\log y)$. Now assume that $1\leqslant d\leqslant t^{2/3}$. Fix $\varepsilon=\frac{1}{10}$. By [40, Ch. III.5, Exercise 293(c)],
$\Psi(x,y):=\#\{n\leqslant x:P^{+}(n)\leqslant y\}\ll xu^{-u}+x^{\varepsilon},\qquad u=\frac{\log x}{\log y},$
uniformly for all $x\geqslant 2$ and $y\geqslant 2$. Let
$v=\frac{\log\frac{t}{d}}{\log y}=s-\frac{\log d}{\log y},$
so that $v\geqslant s/3\geqslant 10$. By partial summation,
[…]
The lower bound $y\geqslant(\log t)^{2}$ implies that $\frac{1+\log s}{\log y}\leqslant\frac{1}{2}$ and thus by the mean value theorem,
$v\log v\geqslant s\log s-(1+\log s)\frac{\log d}{\log y}\geqslant s\log s-\frac{\log d}{2}.$
It follows that the sum of terms on the right side of (3.6) corresponding to $d\leqslant t^{2/3}$ is $\ll(\log y)e^{-s\log s}+t^{\varepsilon-1}$. Since $s\log s\leqslant\frac{1}{2}\log t$ and $s\log(s+1)=s\log s+O(1)$, the right side of (3.6) is
$\ll(\log y)e^{-s\log(s+1)}+(\log y)t^{-2/3}\ll(\log y)e^{-s\log(s+1)},$
as desired. ∎